OpenStudy (mendicant_bias):

(ODE) Dealing with a straightforward IVP, am given initial conditions *and* a family of functions that is the general solution to the IVP, just need to find a member of the family that is a solution as well.

OpenStudy (mendicant_bias):

http://i.imgur.com/rTlwRkY.png

OpenStudy (mendicant_bias):

(Upper bound is infinity, some characters don't display right in that PDF)

ganeshie8 (ganeshie8):

should be straightforward yeah

OpenStudy (mendicant_bias):

Alright, so I have the family of solutions, now I think I just need to like, take derivatives and plug in or something and solve for constants

OpenStudy (mendicant_bias):

Is that right?

OpenStudy (dan815):

ye

OpenStudy (mendicant_bias):

Alright, lol, now the difference between saying it and doing it, going to take a shot now, but in general I'm pretty bumbly/not-good with remembering this stuff in practice. One sec.

OpenStudy (mendicant_bias):

Taking the derivatives:

OpenStudy (mendicant_bias):

\[y=c_{1}e^{x}+c_{2}e^{-x};\]\[y'=c_{1}e^{x}-c_{2}e^{-x}\]

OpenStudy (mendicant_bias):

\[y''=c_{1}e^{x}+c_{2}e^{-x}\]

ganeshie8 (ganeshie8):

y'' is not needed

OpenStudy (mendicant_bias):

Plugging in the constants/conditions:

OpenStudy (mendicant_bias):

Oh, okay, uh, why not?

ganeshie8 (ganeshie8):

use the initial conditions for y and y'

OpenStudy (mendicant_bias):

OH

ganeshie8 (ganeshie8):

you will get two equations solve the constants

OpenStudy (mendicant_bias):

Yeah, I misread, I'm trying to go quick and misread it as an IVP condition

OpenStudy (mendicant_bias):

\[c_{1}e^{x}-c_{2}e^{-x}=y'; \ \ \ y'(0)=c_{1}e^{x}-c_{2}e^{-x}=1.\]\[c_{1}e^{x}=1+c_{2}e^{-x}.\]

OpenStudy (mendicant_bias):

(I'm going about this in the wrong way, I see what I need to do)

OpenStudy (mendicant_bias):

Just add the equations together without rearranging.

ganeshie8 (ganeshie8):

\[y(0) = 0 \implies c_1+c_2 = 0\tag{1}\] \[y'(0) = 1 \implies c_1-c_2 = 1\tag{2}\]

OpenStudy (mendicant_bias):

Oh, because x is zero, yeah, but works the same way, add them together, eliminate c_{2}, 2c_{1}=1, c_{1}= 1/2

OpenStudy (mendicant_bias):

\[2c_{1}=1; \ \ \ c_{1}=\frac{1}{2}\]

OpenStudy (mendicant_bias):

\[\frac{1}{2}-c_{2}=1; \ \ \ c_{2}=-\frac{1}{2}\]

OpenStudy (mendicant_bias):

Lemme check book answers

OpenStudy (mendicant_bias):

Yup, that's right, now moving on to a different type of problem, one minute.